QUESTION FOUR

(5)

(5)

(25 MARKS)

 The following are the variations of ratings of Five Richfield college researchers on the institutions National Innovative Research Initiative (NIRI) project for the year 2020:

RATINGS 3 4 7 10 15 30

Using the data values given above, calculate the following:

4.1.1 Arithmetic mean (5) 4.1.2 Median (5)

   

        4.2 A manufacturer of children toys claims that less than 3% of his products are defective. When 500 toys were drawn from a large production run, 5% were found to be defective.

Based on the above, identify the following:

4.2.1 Population of interest (3) 4.2.2 Sample (3) 4.2.3 Parameter (4) 4.2.4 Statistic (5)


1
Expert's answer
2021-05-14T03:32:02-0400

4.1.1.

"Mean = \frac{3+ 4 +7 +10 +15 +30}{6}=11.5 \\

Median = \frac{7+10}{2}=8.5"

4.2.1 Population of interest - children toys

4.2.2 Sample - 500 toys

4.2.3 Parameter - true value of defective children's toys

4.2.4 Statistic - that 5% of the toys in the sample were found defective.